Use the following for Q1 and Q2. Suppose in a class of 140 students 35 of them are Biology Majors. Q1 Q1. ( point) If a student is chosen at random what is the likelihood they will be a Biology Major? P(Biology Major) Q2. (1 point) If a student is chosen at random what is the likelihood they are not a Biology Major? P(Not Biology Major) Use the following for Q3 and Q5. Suppose at an ice cream shop there are two flavors of ice cream and three different toppings. Assume a person can only choose one type of ice cream; chocolate or vanilla and one topping; sprinkles, nuts or hot fudge. They must have an ice cream with a topping. Q3. (1 point). How many different combinations of ice cream and toppings can be ordered? List the different combinations. This is considered a sample space. 04. (1 point). Assume all combinations are equally likely. What is the probability someone will choose a combination with chocolate ice cream? Q5. ( point). Assume all combinations are equally likely. What is the probability someone will choose a combination with sprinkles?
Answer #1

Solution:-

Given,

Total Students in the class=140

Students with Biology Majors=35

1.)Likelihood they will be a biology major:-

P(Biology Major)=35/140

=0.25

2.)

P(Not Biology Major)=1-P(Biology Major)

=1-0.25

=0.75

I have loved 2 questions here, for rest repost rest of the questions.